# InstructCV: Instruction-Tuned Text-to-Image Diffusion Models as Vision Generalists

### 14 Mar 2024 v2

@@ 4.2 Results
Depth Est. Sem. Seg. Classification Object Det. Denoising Deraining Enhance
@@
SSA [52] $47.2$
M-Former [53] $\mathbf{56.1}$
ResNet [54] $97.5$
ViT [55] $\mathbf{98.4}$
DETR [28] $\mathbf{60.6}$
M-RCNN [56] $60.5$
Uformer [56] $60.5$
MPRNet [56] $60.5$
MIRNet-v2 [56] $60.5$
                      Generalist framework, Task-specific models -
@@
Unified-IO [15] $0.387$ $0.287$ $25.7$ $27.7$ $96.5$ $\mathbf{89.9}$
Pix2SeqV2 [17] 57.4 38.5 - - - - - -
Painter[18] 0.2880.288{0.288}0.288 $0.285$ $49.9$ ${52.3}$
InstructDiff[57] {-}- {-}- {-}- {-}- {-}- {-}- {-}- {-}- - - - - - -
InstructCV-FP $\mathbf{0.275}$ $\mathbf{0.268}$ $\mathbf{52.3}$ $\mathbf{53.5}$ $80.4$ $75.1$ $49.1$ $\mathbf{62.0}$
InstructCV-RP $\mathbf{0.297}$ $\mathbf{0.279}$ $47.2$ $\mathbf{52.1}$ $82.1$ $74.7$ $48.5$ $\mathbf{61.7}$
@@ 5 Conclusion
Limitations. The inference speed of our model lags behind specialized task-specific models and falls short of meeting the real-time inference requirements for tasks such as object detection and segmentation. Additionally, the Pix2Pix formulation can lead to inadmissible output images for a given vision task. Furthermore, the semantic flexibility of InstructCV is constrained by the richness and diversity of our instruction-tuning dataset, which is currently generated by rephrasing a limited set of template prompts. This raises questions for future work: can this learning paradigm accommodate instructions that introduce more nuanced conditions? For example, an instruction might cap the count of objects to be detected. Exploring such ideas might require the integration of strategies such as learning from human feedback, which could enable more versatile generalist models by improving alignment of task outputs with more complex prompts.

### 16 Mar 2024 v3

@@ 4.2 Results
Depth Estimation Semantic Segmentation Classification Object Detection
@@
SSA [52] $47.150$
Mask2Former [53] $\mathbf{56.100}$
ResNet [54] $97.500$
ViT [55] $\mathbf{98.400}$
DETR [28] $\mathbf{60.600}$
Mask R-CNN [56] $60.500$
                  Generalist framework, Task-specific models
@@
Unified-IO [15] $0.387$ $0.287$ $25.713$ $27.724$ $96.514$ $\mathbf{89.877}$
Pix2SeqV2 [17] 57.400 38.500
Painter [18] 0.288 $0.285$ 49.9 52.5
InstructCV-FP $\mathbf{0.275}$ $\mathbf{0.268}$ 52.3 $\mathbf{53.5}$ $80.4$ $75.1$ $49.1$ $\mathbf{62.00}$
InstructCV-RP $0.297$ 0.279 $47.235$ $52.125$ $82.135$ $74.665$ $48.500$ 61.700
@@ 5 Conclusion
Limitations. While InstructCV improves upon the computational cost of existing generalist models, the inference speed of our model lags behind specialized task-specific models and falls short of meeting the real-time inference requirements for tasks such as object detection and segmentation. Additionally, the Pix2Pix formulation can lead to inadmissible output images for a given vision task, e.g., failure to generate a box-shaped output for the object detection task. Furthermore, the semantic flexibility of InstructCV is constrained by the richness and diversity of our instruction-tuning dataset, which is currently generated by rephrasing a limited set of template prompts. This raises questions for future work: can this learning paradigm accommodate instructions that introduce more nuanced conditions? For example, an instruction might cap the count of objects to be detected. Exploring such ideas might require the integration of strategies such as learning from human feedback, which could enable more versatile generalist models by improving alignment of task outputs with more complex prompts.